OpenStudy (anonymous):

MEDAL & FAN ANYBODY ? Find the surface area of the sphere with the given dimension. Leave your answer in therms of pi. Diameter of 16cm. A. 256 cm pi B. 128 cm pi C. 32 cm pi D. 1,024 cm pi

OpenStudy (sleepyjess):

\(4\pi r^2\) is the formula for this one

OpenStudy (anonymous):

i got 803.84 ... @sleepyjess but that doesnt seem right .

OpenStudy (sleepyjess):

remember you are leaving it in terms of pi

OpenStudy (sleepyjess):

do you know what that means?

OpenStudy (anonymous):

@sleepyjess no i dont . Can you help me ?

OpenStudy (sleepyjess):

leaving it in terms of pi means you will do everything except for multiply by \(\pi\)

OpenStudy (sleepyjess):

so \(8^2 ~is~64\\64*4=256~now~add~\pi~to~the~end~and~that~is~your~answer\)

OpenStudy (anonymous):

so its A ? @sleepyjess

OpenStudy (sleepyjess):

\(\Huge\frak\color{lime}{correct}\\\Huge\cal\color{magenta}{you~do~know~that~pi~is~the~same\\~as~\pi~right?}\)

OpenStudy (anonymous):

yes i do :) thanks

OpenStudy (sleepyjess):

ok just making sure \(\LARGE\ddot\smile\)

OpenStudy (sleepyjess):

if you have any other questions you can just tag me in them